OpenStudy (anonymous):

If the midpoint between (x, 6) and (-9, 14) is (8, 10), find the value of x.

OpenStudy (anonymous):

Please

hero (hero):

Hi

OpenStudy (anonymous):

hi Hero. Can you help me please?

hero (hero):

\[(x_{\mid}, y_{\mid})=\left( \frac{x_1 + x_2}{2}, \frac{y_1+ y_2}{2} \right)\]

OpenStudy (anonymous):

Wait nevermind i got the answer already, I put the wrong one..

hero (hero):

I'm sorry to here that

OpenStudy (anonymous):

@Hero here